# MR-STGN: Multi-Residual Spatio Temporal Graph Network using Attention Fusion for Patient Action Assessment

### v1

@@ IV-B Experimental Results
The MR-STGN model was implemented in Python 3.6 using the TensorFlow 2.x framework on a PC with an Intel Xeon Silver 4215R CPU, 32GB of RAM, and a GeForce GTX 3080 Ti 16GB RAM graphics card. The model was trained for 1500 epochs using the Adam optimizer and batch sizes of 4. The learning rate was set to $41e-$ 4. We utilized Huber loss in our study, benefiting from its robustness, balanced nature between MSE and MAE, and reduced sensitivity to outliers. The best model was selected based on its performance on the validation set and was evaluated on the test set. The used evaluation metric is the mean absolute error (MAE) metric.
We present the results obtained from various state-of-the-art approaches, such as the spatio-temporal neural network model, deep CNN model, and deep LSTM model as proposed in [7]. Additionally, we include the co-occurrence model [21], PA-LSTM [22], two-stream CNN [23], and hierarchical LSTM [20]. Within our evaluation, we offer the findings derived from performing an ablation study on our model. This study involved using the same parameters with variations in data input, specifically focusing on positional data alone, angular data alone, and a combination of both.
The results in the table I showcase the performance of various approaches, including our proposed method, across multiple exercises. Our approach, utilizing both positional and angular data, consistently demonstrates superior performance compared to other methods. When considering only positional or angular data, our approach outperforms other individual data-based approaches as well. Moreover, state-of-the-art approaches offer only limited performance for all exercise types presented in the dataset. It is worth mentioning that Exercise 10 (Standing shoulder scaption) is relatively easy to evaluate because it involves a single plane of motion and a few patient body joints. However, it is not perfectly assessed compared to our proposed approach, which provides a good assessment score for this exercise. The average error for our method is the lowest, indicating its effectiveness in accurately assessing rehabilitation exercises. The results from the time execution table II highlight the speed of our model compared to Deb et al. approach [19]. During the training phase, our model for the whole considered 134 videos taking only 29 minutes. Moreover, during the testing phase, our model demonstrated important speed, processing 34 exercise videos in just 3.68 seconds. These results underscore the advantage of our model in terms of both training and testing time, providing a more efficient and practical solution for evaluating rehabilitation exercises.
Fig. 4 demonstrates the computation of joint roles for non-expert users by the attention joint vectors $\chi^{t}$. It reveals that these roles differ from the expert’s pattern when the patient receives a low assessment score. The visualization includes the Deep squat exercise (where the subject bends the knees to descend the body toward the floor with the heels on the floor, the knees aligned over the feet, the upper body remains aligned in the vertical plane) and the Hurdle step exercise (where the subject steps over the hurdle, while the hips, knees, and ankles of the standing leg remain vertical) along with their respective $\chi^{t}$. Refer to Fig. 3 to identify the main joints considered for each exercise.
Exercise Ours with both Only positional data (ours) Only angular data (ours) Deb et al. [19] Du et al. [3] Deep CNN [7] Liao et al. [7] Deep LSTM [7] Co-occurence [21] PA-LSTM [22] Hierarchical LSTM [20] Two-stream CNN [23] Ex1 0.008 0.012 0.011 0.012 0.009 0.013 0.010 0.016 0.010 0.018 0.030 0.288 Ex2 0.010 0.014 0.011 0.011 0.020 0.029 0.028 0.049 0.029 0.044 0.077 0.223 Ex3 0.010 0.016 0.013 0.015 0.036 0.041 0.040 0.093 0.055 0.080 0.137 0.204 Ex4 0.008 0.011 0.010 0.010 0.014 0.016 0.011 0.016 0.013 0.023 0.035 0.360 Ex5 0.007 0.013 0.014 0.010 0.014 0.013 0.018 0.025 0.016 0.031 0.063 0.123 Ex6 0.010 0.015 0.018 0.017 0.020 0.023 0.017 0.021 0.018 0.034 0.046 0.211 Ex7 0.020 0.021 0.025 0.023 0.021 0.033 0.038 0.041 0.027 0.049 0.192 0.050 Ex8 0.020 0.026 0.021 0.024 0.022 0.029 0.023 0.046 0.024 0.050 0.072 0.043 Ex9 0.014 0.017 0.019 0.017 0.025 0.024 0.022 0.044 0.027 0.043 0.065 0.144 Ex10 0.015 0.026 0.028 0.025 0.026 0.036 0.041 0.052 0.046 0.077 0.160 0.110 Average 0,012 0,017 0,017 0,016 0,021 0,026 0,025 0,040 0,027 0,045 0,088 0.175
Phase Number of Videos Execution Time Deb et al. [19] Train 134 22 hours Our model Train 134 21 min Deb et al. [19] Test 34 13.87 seconds Our model Test 34 3.68 seconds
Figure 3: Joints in the skeletal model from UI-PRMD dataset.
@@ V Conclusion
In summary, this paper introduces a novel approach called MR-STGAN, which utilizes 3D skeletons and graph convolutional networks for evaluating patient actions in physical rehabilitation exercises. The proposed model incorporates a multi-residual architecture to capture the spatio-temporal dynamics of patient actions, leveraging information from multiple residual layers. An attention fusion mechanism is employed to selectively combine angular and positional skeleton data, allowing for adaptive weighting of different features and personalized feedback to patients. The performance evaluation on the UI-PRMD dataset demonstrates that the proposed model surpasses existing state-of-the-art models in terms of quantitative measures. Furthermore, the model provides valuable guidance on specific body parts and movements. Its real-time performance makes it suitable for healthcare applications requiring timely feedback. Future work aims to explore continuous assessment possibilities and develop a graphical interface for visual feedback to patients.

### v2

@@ IV-B Experimental Results
The MR-STGN model was implemented in Python 3.6 using the TensorFlow 2.x framework on a PC with an Intel Xeon Silver 4215R CPU, 32GB of RAM, and a GeForce GTX 3080 Ti 16GB RAM graphics card. The model was trained for 1500 epochs using the Adam optimizer and batch sizes of 4. The learning rate was set to $41e-$ 4. The choice of the regression loss function is critical to the performance of a model. Three different regression losses, namely MSE, Log-Cosh, and Huber loss, were used to evaluate the proposed model’s performance. The results show that Huber loss performs better for predicting assessment scores. Huber loss is advantageous as it improves the robustness of the model and provides a balance between MSE and MAE. It is also less sensitive to outliers, making the optimization process more stable and less prone to extreme values in the data. The best model was selected based on its performance on the validation set and was evaluated on the test set. To compare the model’s performance with recent works, 10 runs of the model were performed on the training and testing datasets. The performance measures MAE were saved from each run to ensure accuracy before averaging the results. we present the results obtained from various state-of-the-art approaches, such as the spatio-temporal neural network model, deep CNN model, and deep LSTM model as proposed in [liao2020deep]. Additionally, we include the co-occurrence model [li2018co], PA-LSTM [shahroudy2016ntu], two-stream CNN [simonyan2014two], and hierarchical LSTM [vakanski2018data]. Within our evaluation, we offer the findings derived from performing an ablation study on our model. This study involved using the same parameters with variations in data input, specifically focusing on positional data alone, angular data alone, and a combination of both.
@@
Figure 4: Joints in the skeletal model of Kinect-recorded data from UI-PRMD dataset.
@@ V Conclusion
In conclusion, this paper proposes a novel Multi-Residual Spatio Temporal Graph Network (MR-STGAN) that utilizes 3D skeletons and graph convolutional networks to assess patient actions during physical rehabilitation exercises. The proposed model employs a multi-residual architecture that captures the spatio-temporal dynamics of patient actions by integrating information from multiple residual layers. An attention fusion mechanism is utilized to selectively combine the information from the angular and positional skeletons, allowing the network to adaptively weight the importance of different features and provide more accurate and personalized feedback to patients. The performance of the proposed model was evaluated on physical rehabilitation dataset UI-PRMD, and it was found to outperform existing state-of-the-art models in terms of quantitative measures. Additionally, the proposed model provides clear guidance on which body parts or movements to focus on, and it has shown high performance on physical rehabilitation dataset. The real-time performance of the model makes it suitable for healthcare applications requiring timely feedback. In future work, the authors intend to explore the possibility of continuous assessment and develop a graphical interface to provide visual feedback to patients.